OpenStudy (anonymous):

Is this equation correct: (x y)3 = x y3 ? Explain why. If it is correct, how can you prove this? If it is incorrect, what should be done to make it right?

OpenStudy (anonymous):

loco85 Is this equation correct: (x y)3 = x y3 ? Explain why. If it is correct, how can you prove this? If it is incorrect, what should be done to make it right?

OpenStudy (anonymous):

amistre64 Is this equation correct: (x y)3 = x y3 ? Explain why. If it is correct, how can you prove this? If it is incorrect, what should be done to make it right?

OpenStudy (anonymous):

I ssume you mean \[(xy)^3 = xy^3\] no, this is not correct, you must cube both terms, thus: \[(xy)^3 = x^3timesy^3\]

OpenStudy (anonymous):

thanks guyc i appreciate it